aki's bicycle designs has determined that when x hundred bicycles are built, the average cost per bicycle is given by C(x)=0.5x^2-1.2x+2.725, how many bikes should the shop build to minimize the average cost per bike? thank you
Answer by mathbath(13)

The given function to determine the cost per bicycle is ..
C(x)=0.5x^2-1.2x+2.725 .
The standard form of a quadratic equation is ..
f(x)=ax^2+bx+c
Comparing the above two equations we can say that the Cost function C(x) is in a standard quadratic equation form where..
a = 0.5, b = -1.2, c = 2.725
Now if any quadratic equation is plotted on a graph we get a parabola (which looks like the letter "U" or an inverted "U")
For quadratic equations..
if a > 0 , The parabola is in the form of a letter "U" i.e., it opens upwards. This also means that the function has a minimum which is at the vertex of the parabola.
Considering the standard form of quadratic equation, the vertex of the parabola has co-ordinates {-b/2a , f(-b/2a)}
So,
x = -b/2a
Substituting the values for a and b we can find the value of x.
x = -(-1.2)/2(0.5) = 1.2/1 = 1.2
So, the Cost function has a minimum when x = 1.2.
Since the Cost function is for 100x bicycles , aki's bicycles needs to manufacture atleast 100(1.2) bicycles to minimize the cost per bicycle.
So the total no. of required bicycles is 120
Verify:
Now lets verify our answer.
The value of x that we got is 1.2. Lets substitute this in the Cost function C(x)
C(x) = 0.5(1.2)^2 - 1.2(1.2) + 2.725
= 0.5(1.44) - 1.44 + 2.725
= 0.72 + 1.285
= 2.005
Lets raise the value of x to 2 and substitute in the cost function C(x)
C(x) = 0.5(2)^2 - 1.2(2) + 2.725
= 0.5(4) - 2.4 + 2.725
= 2 + 0.325
= 2.325
So as you can see as we raised the value of x from 1.2 to 2 the C(x) function jumped up from 2.005 to 2.325. In other words what this means is the value of x=1.2 is the minimum for this function.
Hope this helps.